OpenStudy (anonymous):

For x not equal to 0, the slope of the tangent to y=xcosx equals zero whenever 1.) tanx=-x 2.)tanx=(1/x) 3.)tanx=x 4.)sinx=x 5.)cosx=x Please help, explanation would be greatly appreciated.

OpenStudy (anonymous):

Sometimes, you just gotta learn to listen to those whispers in your ear ^^

OpenStudy (anonymous):

So, when you hear "slope of the tangent", something should always whisper "derivative" in your ear :D What's the derivative of xcos(x) ?

OpenStudy (anonymous):

x-sinx+cosx

OpenStudy (anonymous):

Umm... somehow, I feel you have the right idea, but your answer is off ^^ There should only be two terms.

OpenStudy (anonymous):

I bet you meant x(-sin(x)) + cos(x)

OpenStudy (anonymous):

Yes! I did I'm sorry

OpenStudy (anonymous):

Be more careful next time ^^ Your teacher will probably not be as considerate :D Anyway, this comes down to y' = cos(x) - xsin(x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, this derivative IS the slope of the tangent line. So... when is this equal to zero? 0 = cos(x) - xsin(x)

OpenStudy (anonymous):

I'm not sure to be honest

OpenStudy (anonymous):

I'm not sure how to find that

OpenStudy (anonymous):

You just have to be creative ^^ Here. It means \[\large \cos(x) = x\sin(x)\] Maybe you can work it out from here

OpenStudy (anonymous):

oh shoot

OpenStudy (anonymous):

x=cosx/sinx

OpenStudy (anonymous):

and cosx/sinx=tanx

OpenStudy (anonymous):

uhhh, no. \[\large \frac{\sin(x)}{\cos(x)}=\tan(x)\]

OpenStudy (anonymous):

wow sorry as you can tell I'm really bad at this

OpenStudy (anonymous):

But this should put you on the right track ^^

OpenStudy (anonymous):

Sure, you're right about \[\large x = \frac{\cos(x)}{\sin(x)}\] This should bring you really close to the answer :D

OpenStudy (anonymous):

would it be 1/x? becasue we want to change sinx/cosx into cosx/sinx

OpenStudy (anonymous):

That's right ^^

OpenStudy (anonymous):

I understand now, i just have to remember cosx/sinx is equal to tanx. thank you

OpenStudy (anonymous):

As you can see

OpenStudy (anonymous):

You're not as bad at this as you think :)

OpenStudy (anonymous):

:) thank you